Consider a market described by the linear demand curve q = 8 − 2p.
(e) Compute the price elasticity of demand at the prices p = 3, p = 2, and p = 1.

At which of these prices is demand elastic / inelastic / unit-elastic?
Answer #1

Answer

да дp Q Price elasticity of demand = ana Where, да Ә(8 – 2р) - 2 др др

Now, doing one by one for the prices provided

p=3

q=8-2*3=2

Price elasticity of demand = -2* 2000 do

The demand is elastic as the elasticity is below -1.

=========

p=2, q=8-2*2=4

Price elasticity of demand = -2* AN

the demand is unit elastic as the elasticity is equal to -1.

==========

p=1,9 = 8 – 2*1 = 6 Price elasticity of demand = -2*== = -0.33

the demand is inelastic as the elasticity is above -1.